# Exceptional supercuspidal representations in small residue characteristic

### v1

@@ 4.3 $G=G_{2}$, $p=3$
When $G=G_{2}$, by [FR17, Theorem 4.2], the points $y\in\mathcal{A}$ at which $V_{y}^{\vee}$ admits stable vectors are conjugate under the affine Weyl group to $\frac{1}{m}(\omega_{1}^{\vee}+\omega_{2}^{\vee})$ for $m=2,3,6$. Under the construction of [RY14], these stable vectors give rise to epipelagic representations of depth $1/2,1/3$ and $1/6$ respectively, by Proposition 3.10. The point $y\in\mathcal{A}$ with Kac coordinates $01\Rrightarrow 0$ is conjugate to $\frac{1}{2}(\omega_{1}^{\vee}+\omega_{2}^{\vee})$. As $\pi_{\lambda}$ has depth $1/2$, it suffices to show that $\pi_{\lambda}$ is not one of the representations $\pi_{y}(\lambda^{\prime},\rho)$ where $\lambda^{\prime}\in V_{y}^{\vee}(\mathbb{F}_{q})$ is a stable vector and $\rho\in\mathrm{Irr}(\mathcal{H}_{\lambda^{\prime}})$, as defined in Section 3.2.
@@
Recall (for example, [GR10, Section 7]) that the formal degree with respect to $\mu$ of a compactly induced representation $\mathrm{ind}_{J}^{G(k)}V$ is $\dim V/\mu(J)$. By Proposition 3.7, if $\pi_{\lambda}$ is isomorphic to some $\pi(\lambda^{\prime},\rho)$, then we must have
@@
$\left(\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right),\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right),\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)$. (5)
@@
where $a,b,c,d,e,f,g,h\in\overline{\mathbb{F}}_{3}$. If $F$ is stabilised by one of the elements of (5), one checks that $d=f=0$ or $e=f=0$. By [Rom16, Proposition 2.15], there is a polynomial $\Delta(F)$ in $a,\dots,h$ such that $F$ is stable for the action of $\mathrm{SL}_{2}(\overline{\mathbb{F}}_{3})\times\mathrm{SL}_{2}(\overline{\mathbb{F}}_{3})$ if and only if $\Delta(F)\neq 0$. Explicitly, there are polynomials $H_{6}(F)$ and $G_{6}(F)$ such that $\Delta(F)=H_{6}(F)^{3}G_{6}(F)$ and $H_{6}(F)$ reduces in characteristic 3 to $c^{3}f^{3}-d^{3}e^{3}$. We see that this is 0 for any $F$ stabilised by an element of order 3, so that such $F$ is not a stable vector.
Remark 4.12.

### v2

@@ 4.3 $G=G_{2}$, $p=3$
When $G=G_{2}$, by [FR17, Theorem 4.2], the points $y\in\mathcal{A}$ at which $V_{y}^{\vee}$ admits stable vectors are conjugate under the affine Weyl group to $\frac{1}{m}(\omega_{1}^{\vee}+\omega_{2}^{\vee})$ for $m=2,3,6$. Under the construction of [RY14], these stable vectors give rise to epipelagic representations of depth $1/2,1/3$ and $1/6$ respectively, by Proposition 3.2.6. The point $y\in\mathcal{A}$ with Kac coordinates $01\Rrightarrow 0$ is conjugate to $\frac{1}{2}(\omega_{1}^{\vee}+\omega_{2}^{\vee})$. As $\pi_{\lambda}$ has depth $1/2$, it suffices to show that $\pi_{\lambda}$ is not one of the representations $\pi_{y}(\lambda^{\prime},\rho)$ where $\lambda^{\prime}\in V_{y}^{\vee}(\mathbb{F}_{q})$ is a stable vector and $\rho\in\mathrm{Irr}(\mathcal{H}_{\lambda^{\prime}})$, as defined in Section 3.2.
@@
Recall (for example, [GR10, Section 7]) that the formal degree with respect to $\mu$ of a compactly induced representation $\mathrm{ind}_{J}^{G(k)}V$ is $\dim V/\mu(J)$. By Proposition 3.2.3, if $\pi_{\lambda}$ is isomorphic to some $\pi(\lambda^{\prime},\rho)$, then we must have
@@
$\left(\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right),\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right),\left(\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\begin{pmatrix}1&1\\ 0&1\end{pmatrix}\right)$. (4.3.7)
@@
where $a,b,c,d,e,f,g,h\in\overline{\mathbb{F}}_{3}$. If $F$ is stabilised by one of the elements of (4.3.7), one checks that $d=f=0$ or $e=f=0$. By [Rom16, Proposition 2.15], there is a polynomial $\Delta(F)$ in $a,\dots,h$ such that $F$ is stable for the action of $\mathrm{SL}_{2}(\overline{\mathbb{F}}_{3})\times\mathrm{SL}_{2}(\overline{\mathbb{F}}_{3})$ if and only if $\Delta(F)\neq 0$. Explicitly, there are polynomials $H_{6}(F)$ and $G_{6}(F)$ such that $\Delta(F)=H_{6}(F)^{3}G_{6}(F)$ and $H_{6}(F)$ reduces in characteristic 3 to $c^{3}f^{3}-d^{3}e^{3}$. We see that this is 0 for any $F$ stabilised by an element of order 3, so that such $F$ is not a stable vector.
Remark 4.3.8.